OpenStudy (anonymous):

( a vector+ B vector) . (a vector- B vector)= 2A squared + AB cos theta- 6B squared

OpenStudy (anonymous):

Prove the above

OpenStudy (anonymous):

Define x axis to be on A. \[X \cdot Y=X_xY_x+X_yY_y\]

OpenStudy (anonymous):

\[(A+B) \cdot(A-B)= (A+B)_x (A-B)_x+(A+B)_y (A-B)_y\]A has no y component, so this reduces to \[ (A+B)_x (A-B)_x+(B)_y (-B)_y \] And I think \[ (A+B)_x =A_x +B_x \]

OpenStudy (anonymous):

Hey i've not quite comprehended that